the speed of a freight train is 12km/h slower than the speed of a passenger train the freight train travel 348km in the same time it takes the passenger train to travel 420 km find the speed of the freight train
Answer by Cromlix(4381)

Passenger train = x
Freight train = x - 12
348km/x-12 = 420/x
Cross multiply
348x = 420(x - 12)
348x = 420x - 5040
-72x = -5040
-x = -70
x = 70km/hr
Passenger train's speed = 70 km/hr
Freight train's speed = 58 km/hr
Proof:
348/58 = 420/70
6 = 6
Hope this helps.
:-)